A solenoidal coil with 23 turns of wire is wound tightly around another coil with 320 turns. The inner solenoid is 25.0 cm long and has a diameter of 2.30 cm . At a certain time, the current in the inner solenoid is 0.100 A and is increasing at a rate of 1700 A/s



Calculate the emf induced in the outer solenoid by the changing current in the inner solenoid
Answer #1

Plz upvote
Answer #2

Given:

  • Turns of outer solenoid, N2=23

  • Turns of inner solenoid, N1=320

  • Length of inner solenoid, l=25.0cm=0.25m

  • Diameter of inner solenoid, d=2.30cmr=1.15×102m

  • Current in inner solenoid, I=0.100A

  • Rate of current change, dIdt=1700A/s

Step 1: Calculate mutual inductance (M)
The mutual inductance between two tightly coupled solenoids is:

M=μ0N1N2Al

where , and A=πr2.

A=π(1.15×102)2=4.154×104m2M=(4π×107)(320)(23)(4.154×104)0.25M1.53×105H(15.3 μH)

Step 2: Calculate induced emf in outer solenoid
Faraday's Law:

E=MdIdtE=(1.53×105)(1700)=0.0260V

(Magnitude: 0.0260V, sign indicates direction.)


Answer:
The induced emf in the outer solenoid is 0.0260V


answered by: anonymous